# Field-based Molecule Generation

### v1

@@ A.2 Conditional generation
A.3 Chirality Analysis
Experiment
@@ A.3 Chirality Analysis
Theoretical analysis
Let $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{3}$ be two, three-dimensional vectors, and let $\mathbf{d}=\mathbf{v}_{1}-\mathbf{v}_{2}$. By definition, orthogonal matrices $Q$ have an inverse equal to $Q^{-1}=Q^{T}$. Therefore, it is easy to see that $||Q\mathbf{d}||=||\mathbf{d}||$ since
$||Q\mathbf{d}||^{2}=(Q\mathbf{d})^{T}(Q\mathbf{d})=\mathbf{d}^{T}Q^{T}Q\mathbf{d}=\mathbf{d}^{T}\mathbf{I}\mathbf{d}=\mathbf{d}^{T}\mathbf{d}=||\mathbf{d}||^{2},$ (15)
where $\mathbf{I}$ is the identity matrix, and |||||$\cdot$||| | ⋅ | | is the L2 norm. Since reflection and rotation matrices are orthogonal, it follows, that the Euclidean distances between any two vectors $\mathbf{v}_{1},\mathbf{v}_{2}$ are invariant to orthogonal matrix multiplication, and, therefore, to any rotation or reflection matrix.
$\cos(Q\mathbf{v}_{1},Q\mathbf{v}_{2})=\frac{Q\mathbf{v}_{1}\cdot Q\mathbf{v}_{2}}{||Q\mathbf{v}_{1}||||Q\mathbf{v}_{2}||}=\frac{(Q\mathbf{v}_{1})^{T}(Q\mathbf{v}_{2})}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\frac{\mathbf{v}_{1}\mathbf{I}\mathbf{v}_{2}}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\cos(\mathbf{v}_{1},\mathbf{v}_{2}),$ (16)
where we have used equation 15 for $||Q\mathbf{v}_{1}||=||\mathbf{v}_{1}||$. All distances and angles between all vectors in the $\mathbb{R}^{3}$ space are therefore preserved.
Figure 8: R and S configurations of a general tetrahedral configuration with three, distinct residues.
Consider some tetrahedral configuration, where the central carbon atom is positioned at $\mathbf{m}_{C}$, and is connected to four unique residues $R1,R2,R3$, and $R4$, initially positioned at $\mathbf{m}_{R1},\mathbf{m}_{R2},\mathbf{m}_{R3}$, and $\mathbf{m}_{R4}$, respectively (see Figure 8). Observing the molecule along the $R_{1}$-C-axis, a way of defining the chirality of the molecule is to specify the ordering of the remaining residues in clock-wise or counter-clockwise orientation. Assuming we start with $R2$ position fixed at $\mathbf{m}_{R2}$, swapping $R3$ and $R4$ positions has the effect of reversing the rotation direction of $R3$ and $R4$, and creating a distinct chiral configuration. Now, considering the plane defined by $m_{R1}$, the central carbon position $\mathbf{m}_{C}$, and $\mathbf{m}_{R2}$, the same effect can be achieved by a reflection about this plane.
To our knowledge, all current point-cloud-based methods use Euclidean distances or relative angles between atom position vectors $\mathbf{m}_{i},\mathbf{m}_{j}$, e.g., for message passing in their GNN architectures. These methods cannot distinguish enantiomers, and, as such, will always produce them completely at random, and the molecular graphs will not change based on whether the S or R enantiomer configuration is present in that molecule.
Appendix B Field parameters

### v2

@@ A.2 Conditional generation
Appendix B Chirality Analysis and E(3) invariance
In Appendix B.1, we develop our theoretical analysis and in Appendix B.2 we describe our chirality experiment details.
B.1 Theoretical analysis
In Appendix B.1.1, we draw the connection between a NN’s E(3) invariance and its resulting invariance to enantiomer pairs.
In Appendix B.1.2, we show how point-cloud Diffusion models parametrized by E(3) invariant NN become E(3) equivariant.
We give proofs of Proposition1, Lemma 2, and Proposition 3, in Appendix B.1.3, B.1.4, and B.1.5. In Appendix B.1.6 we discuss cases where Proposition 3 applies.
B.1.1 E(3) invariance of NNs and chirality
Let $\mathbf{v}_{1},\mathbf{v}_{2}\in\mathbb{R}^{3}$ be two, three-dimensional vectors, and let $\mathbf{d}=\mathbf{v}_{1}-\mathbf{v}_{2}$. By definition, orthogonal matrices ${\bm{Q}}$ have an inverse equal to ${\bm{Q}}^{-1}={\bm{Q}}^{T}$. Therefore, it is easy to see that $||{\bm{Q}}\mathbf{d}||=||\mathbf{d}||$ since
$||{\bm{Q}}\mathbf{d}||^{2}=({\bm{Q}}\mathbf{d})^{T}({\bm{Q}}\mathbf{d})=\mathbf{d}^{T}{\bm{Q}}^{T}{\bm{Q}}\mathbf{d}=\mathbf{d}^{T}{\bm{I}}\mathbf{d}=\mathbf{d}^{T}\mathbf{d}=||\mathbf{d}||^{2},$ (15)
where ${\bm{I}}$ is the identity matrix, and |||||$\cdot$||| | ⋅ | | is the L2 norm. Since reflection and rotation matrices are orthogonal, it follows, that the Euclidean distances between any two vectors $\mathbf{v}_{1},\mathbf{v}_{2}$ are invariant to orthogonal matrix multiplication, and, therefore, to any rotation or reflection matrix.
$\cos({\bm{Q}}\mathbf{v}_{1},{\bm{Q}}\mathbf{v}_{2})=\frac{{\bm{Q}}\mathbf{v}_{1}\cdot{\bm{Q}}\mathbf{v}_{2}}{||{\bm{Q}}\mathbf{v}_{1}||||{\bm{Q}}\mathbf{v}_{2}||}=\frac{({\bm{Q}}\mathbf{v}_{1})^{T}({\bm{Q}}\mathbf{v}_{2})}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\frac{\mathbf{v}_{1}^{T}{\bm{I}}\mathbf{v}_{2}}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{||\mathbf{v}_{1}||||\mathbf{v}_{2}||}=\cos(\mathbf{v}_{1},\mathbf{v}_{2}),$ (16)
where we have used equation 15 for $||{\bm{Q}}\mathbf{v}_{i}||=||\mathbf{v}_{i}||$. All distances and angles between all vectors in the $\mathbb{R}^{3}$ space are therefore preserved.
Figure 8: R and S configurations of a general tetrahedral configuration with four, distinct residues (the 4th atom, perpendicular to the planar perspective was not annotated for simplicity). The determinant $\det[\mathbf{m}_{R1}-\mathbf{m}_{C},\mathbf{m}_{R2}-\mathbf{m}_{C},\mathbf{m}_{R3}-\mathbf{m}_{C}]$ is constant under rotations, but changes sign after reflections. Equivalently, the sign changes when swapping $\mathbf{m}_{R1}$ and $\mathbf{m}_{R3}$ positions, which defines the clockwise (left) or counterclockwise (right) orientation. The determinant sign or, equivalently, the orientation determines the enantiomer R or S configuration.
Consider a tetrahedral configuration, where the central carbon atom is positioned at $\mathbf{m}_{C}$, and is connected to four unique residues $R1,R2,R3$, and $R4$, initially positioned at $\mathbf{m}_{R1},\mathbf{m}_{R2},\mathbf{m}_{R3}$, and $\mathbf{m}_{R4}$, respectively (see Figure 8). Observing the molecule along the $R_{4}$-C-axis, a way of defining the chirality of the molecule is to specify the ordering of the remaining residues in clockwise or counterclockwise orientation. Assuming we start with $R2$ position fixed at $\mathbf{m}_{R2}$, swapping $R1$ and $R3$ positions has the effect of reversing the rotation direction and creating a distinct chiral configuration. Now, considering the plane defined by $\mathbf{m}_{R2}$, the central carbon position $\mathbf{m}_{C}$, and $\mathbf{m}_{R4}$, the same effect can be achieved by a reflection about this plane.
By Householder transformation, a reflection about a hyperplane (in this case, plane), can be defined by the plane’s normal vector (perpendicular to the plane and of norm 1), for any plane that passes the origin:
${\bm{P}}={\bm{I}}-2\mathbf{v}\mathbf{v}^{T},$ (17)
where the reflection matrix ${\bm{P}}$ leaves any vector on the plane unchanged. Translating the object by, e.g., $-(\mathbf{m}_{R2}+\mathbf{m}_{C}$ + $\mathbf{m}_{R4}$)/3, the plane now passes the origin, meaning we can determine such a matrix ${\bm{P}}$ for the $\mathbf{m}_{R2},\mathbf{m}_{C}$, $\mathbf{m}_{R4}$, and swap $\mathbf{m}_{R_{3}}$ and $\mathbf{m}_{R_{1}}$ with that reflection.
To our knowledge, all current point-cloud-based methods use Euclidean distances or relative angles between atom position vectors $\mathbf{m}_{i},\mathbf{m}_{j}$, e.g., for message passing in their GNN architectures. These methods cannot distinguish enantiomers. As such, their resulting E(3) invariant probability distributions will always produce them completely at random, and the molecular graphs will not change based on whether the S or R enantiomer configuration is present in that molecule.
B.1.2 E(3) invariance of the generative setting (DDPM)
The problem arising from E(3) invariant neural networks in molecule generation is the E(3) equivariance of the resulting conditional probability distribution $p_{\phi}(\mathbf{m}_{s}|\mathbf{m}_{t})$. Concretely, for a set of positions $\mathbf{m}=\{\mathbf{m}_{1},\ldots,\mathbf{m}_{N}\},\mathbf{m}_{i}\in\mathbb{R}^{3}$, the R and S enantiomer configurations of molecules will be equally likely, since the marginal becomes E(3) invariant $p_{\phi}(\mathbf{m})=p_{\phi}(R\mathbf{m}),\forall R$ orthogonal. We illustrate this in the context of EDM (Hoogeboom et al., 2022). Consider a set of $\mathbb{R}^{3}$ points $\{\mathbf{m}_{1},\ldots,\mathbf{m}_{N}\}$ of a molecule containing $N$ atoms and a neural network $\phi_{\mathbf{m}}$. At each layer $l$ in the GNN network parametrizing EDM, positions are updated as follows:
$\mathbf{m}_{i}^{l+1}=\mathbf{m}_{i}^{l}+\sum_{i\neq j}\frac{\mathbf{m}_{i}^{l}-\mathbf{m}_{j}^{l}}{d_{ij}+1}\phi_{\mathbf{m}}(d_{ij},\ldots),$ (18)
where $\phi_{\mathbf{m}}$ receives atom pair Euclidean distances and other features that are E(3) invariant.
Every layer is a linear combination of atom positions $\mathbf{m}_{i}$ with E(3) invariant coefficients that are equal to 1 or have the form {±$\{\pm\phi_{\mathbf{m}}(d_{ij},\ldots)/(1+d_{ij})\}$. The update at layer $l+1$ of atom $i$ is then given by $\mathbf{m}_{i}^{l+1}=T^{l+1}_{i}(\mathbf{m}^{l})$, and since $T^{l+1}_{i}$’s coefficients are E(3) invariant, the equality $T^{l+1}_{i}({\bm{R}}\mathbf{m}^{l})={\bm{R}}T^{l+1}_{i}(\mathbf{m}^{l})$ holds for all orthogonal matrices ${\bm{R}}$.
Consider now the $\mathbf{\mathbf{\epsilon}}_{\phi}$ parametrization of the reverse DDPM process
$p_{\phi}(\mathbf{m}_{s}|\mathbf{m}_{t},t)=\mathcal{N}(\mathbf{m}_{s};\boldsymbol{\mu}_{\phi}(\mathbf{m}_{s},s,t),\mathbf{\Sigma}(s,t)),$ (19)
where
$\boldsymbol{\mu}_{\phi}(\mathbf{m}_{t};s,t)=\frac{1}{\alpha_{ts}}\mathbf{m}_{t}-\frac{\sigma_{ts}^{2}}{\alpha_{ts}\sigma_{t}}\mathbf{\epsilon}_{\phi}(\mathbf{m}_{t},t),$ (20)
where $\alpha_{ts}=\alpha_{t}/\alpha_{s}$ and $\sigma_{ts}^{2}=\sigma_{t}^{2}-\alpha_{ts}^{2}\sigma_{s}^{2}$, keeping the notation from Section 4.2. It follows that
$\boldsymbol{\mu}_{\phi}({\bm{R}}\mathbf{m}_{t};s,t)=$ $\frac{1}{\alpha_{ts}}{\bm{R}}\mathbf{m}_{t}-\frac{\sigma_{ts}^{2}}{\alpha_{ts}\sigma_{t}}\mathbf{\epsilon}_{\phi}({\bm{R}}\mathbf{m}_{t},t)$
$=$ $\frac{1}{\alpha_{ts}}{\bm{R}}\mathbf{m}_{t}-\frac{\sigma_{ts}^{2}}{\alpha_{ts}\sigma_{t}}{\bm{R}}\mathbf{\epsilon}_{\phi}(\mathbf{m}_{t},t)$ equivariance of $\epsilon_{\phi}$
$=$ ${\bm{R}}(\frac{1}{\alpha_{ts}}\mathbf{m}_{t}-\frac{\sigma_{ts}^{2}}{\alpha_{ts}\sigma_{t}}\mathbf{\epsilon}_{\phi}(\mathbf{m}_{t},t))$
$=$ ${\bm{R}}\boldsymbol{\mu}_{\phi}(\mathbf{m}_{t};s,t),$
Since the covariance matrix of the reverse process $\mathbf{\Sigma}$ does not depend on positions $\mathbf{m}$, and because of the $\boldsymbol{\mu}_{\phi}$ equivariance wrt E(3), we have that
$p_{\phi}(\mathbf{m}_{s}|\mathbf{m}_{t},t)=p_{\phi}({\bm{R}}\mathbf{m}_{s}|{\bm{R}}\mathbf{m}_{t},t),\forall t.$ (21)
We give a short version of the proof from (Xu et al., 2022; Hoogeboom et al., 2022), which shows how the conditional $p_{\phi}(\mathbf{m}_{s}|\mathbf{m}_{t},t)$ and marginal $p_{\phi}(\mathbf{m}_{1})$ being, respectively, equivariant and invariant to a symmetry group, results in the invariance of $p_{\phi}(\mathbf{m}_{0})$. Following (Hoogeboom et al., 2022), we have
$p({\bm{R}}\mathbf{m}_{s})=$ $\int_{\mathbf{m}_{t}}p({\bm{R}}\mathbf{m}_{s}|\mathbf{m}_{t})p(\mathbf{m}_{t})$
$=$ $\int_{\mathbf{m}_{t}}p({\bm{R}}\mathbf{m}_{s}|{\bm{R}}{\bm{R}}^{-1}\mathbf{m}_{t})p({\bm{R}}{\bm{R}}^{-1}\mathbf{m}_{t})$
$=$ $\int_{\mathbf{m}_{t}}p(\mathbf{m}_{s}|{\bm{R}}^{-1}\mathbf{m}_{t})p({\bm{R}}^{-1}\mathbf{m}_{t})$ Invariance of $p(\mathbf{m}_{t})$ equivariance of $p(\mathbf{m}_{s}|\mathbf{m}_{t})$
$=$ $\int_{\mathbf{m}^{\prime}}p(\mathbf{m}_{s}|\mathbf{m}^{\prime})p(\mathbf{m}^{\prime})\cdot|\det{\bm{R}}|$ Change of variables $\mathbf{m}^{\prime}={\bm{R}}\mathbf{m}_{t}$
$=$ $p(\mathbf{m}_{s}),$
where we note that the change of variables $\mathbf{m}^{\prime}={\bm{R}}\mathbf{m}_{t}$ also changes the integration volume, and the equality stands for both $\det{\bm{R}}=\pm 1$ (rotation and reflection matrices).
Observing that $p(\mathbf{m}_{1})=\mathcal{N}(\mathbf{0},\mathbf{I})$ is invariant to orthogonal matrix multiplication, it follows that $p(\mathbf{m}_{0})=p({\bm{R}}\mathbf{m}_{0})$ for all orthogonal matrices ${\bm{R}}$ (induction). We have shown how reflection matrices ${\bm{R}}$ switch the R and S chiral configurations of a molecule at the beginning of this section. This result can be empirically observed in the orange line of the main text Figure 3.
B.1.3 Proposition 1 proof.
[…]
$p_{\phi}(\mathbf{m})=p_{\phi}({\bm{R}}_{+}\mathbf{m}+\mathbf{t}),$ (22)
for all rotation matrices ${\bm{R}}_{+}$ (orthogonal, with $\det{\bm{R}}_{+}=1$) and $\mathbf{t}$ is a translation vector. The E(3) invariance of $p_{\phi}$ gives
$p_{\phi}({\bm{R}}_{-}({\bm{R}}_{+}\mathbf{m}+\mathbf{t}))=p_{\phi}({\bm{R}}_{+}\mathbf{m}+\mathbf{t})=p_{\phi}(\mathbf{m}),$ (23)
where ${\bm{R}}_{-}$ is orthogonal, with $\det{\bm{R}}_{-}=-1$.
We restate that swapping two residue positions $\mathbf{m}_{R1}$ and $\mathbf{m}_{R3}$ of atoms $R1$ and $R3$ connected to a chiral center will change the clockwise or counterclockwise orientation (see Figure 8). Swapping $\mathbf{m}_{R1}$ and $\mathbf{m}_{R3}$ positions can be achieved through a reflection about the plane defined by the points $\{\mathbf{m}_{R4},\mathbf{m}_{C},\mathbf{m}_{R2}\}$. In turn, a reflection about a plane can be expressed as a rotation, a translation, and a reflection. This implies that $\exists{\bm{R}}_{+},{\bm{R}}_{-},\mathbf{t}$, s.t. $\mathbf{m}^{\prime}={\bm{R}}_{-}({\bm{R}}_{+}\mathbf{m}+\textbf{t})$, which gives
$p_{\phi}(\mathbf{m}^{\prime})=p_{\phi}({\bm{R}}_{-}({\bm{R}}_{+}\mathbf{m}+\textbf{t}))=p_{\phi}(\mathbf{m}).$ (24)
B.1.4 Lemma 2 proof.
Let $M^{\prime}=\{\mathbf{m}_{1},\ldots,\mathbf{m}_{m}\}$, $m\leq n,\mathbf{m}_{i}\in\mathbb{R}^{n}$ be $n$-dimensional vectors, and $f:\mathbb{R}^{m\times n}\rightarrow S$ be an SE(n) invariant function. Center the object in 0, creating
$M=\{\mathbf{m}_{1}-\bar{\mathbf{m}},\ldots,\mathbf{m}_{m}-\bar{\mathbf{m}}\}$ (25)
where $\bar{\mathbf{m}}=\frac{1}{m}\sum_{i=1}^{m}\mathbf{m}_{i}$. Observe how the vectors of $M$ are linearly dependent, since
$\sum_{i=1}^{m}(\mathbf{m}_{i}-\bar{\mathbf{m}})=\sum_{i=1}^{m}(\mathbf{m}_{i})-n\cdot\frac{1}{n}\sum_{i=1}^{m}\mathbf{m}_{i}=\boldsymbol{0}$ (26)
This means that $dim(M)\leq m-1\leq n-1$, which gives $dim(M^{\perp})\geq 1$, since $M$ contains at most $n-1$ linearly independent vectors. This gives
$M^{\perp}=\{\mathbf{x}\in\mathbb{R}^{n}|\langle\mathbf{x},\mathbf{m}\rangle=0,\forall\mathbf{m}\in Span(M)\}\neq\{\boldsymbol{0}\}.$ (27)
Let $\mathbf{p}\in M^{\perp}$ be any of $M$’s hyperplane normal vectors (perpendicular to all vectors in $M$ and unit length). Using the Householder transformation, define a reflection matrix about the hyperplane using
${\bm{P}}={\bm{I}}-2\mathbf{p}\mathbf{p}^{T}.$ (28)
Note how any vector in $M$ is unchanged when reflecting by $\mathbf{P}$, since
${\bm{P}}\mathbf{m}=({\bm{I}}-2\mathbf{p}\mathbf{p}^{T})\mathbf{m}=\mathbf{m}-2\mathbf{p}\mathbf{p}^{T}\mathbf{m}=\mathbf{m}-\boldsymbol{0},\forall\mathbf{m}\in M.$ (29)
For all ${\bm{R}}$, $det({\bm{R}})=-1$ and ${\bm{R}}$ orthogonal (reflection matrix), note how
[…]
$=f({\bm{R}}M)$ since $\text{since }{\bm{P}}M=M$ (31)
where ${\bm{R}}{\bm{P}}$ is a proper rotation matrix since it is orthogonal and $det({\bm{R}}{\bm{P}})=det({\bm{R}})det({\bm{P}})=-1\cdot-1=1$.
We have established that $f$ is reflection invariant for 0-centered objects of $m\leq n$ $\mathbb{R}^{n}$ vectors. We will utilize the translation invariance of $f$ to extend the proof to any $M^{\prime}$ object of $m\leq n$ $\mathbb{R}^{n}$ vectors.
Let $T_{\mathbf{v}}$ be a translation operator applied to a set of vectors:
$T_{\mathbf{v}}(M^{\prime})=\{\mathbf{m}_{1}+\mathbf{v},\ldots,\mathbf{m}_{m}+\mathbf{v}\}.$ (32)
Let ${\bm{R}}$ be a reflection matrix. Then
[…]
$=f({\bm{R}}\{\mathbf{m}_{1}-\bar{\mathbf{m}}+\bar{\mathbf{m}},\ldots,\mathbf{m}_{m}-\bar{\mathbf{m}}+\bar{\mathbf{m}}\})$
$=f(\{{\bm{R}}\mathbf{m}_{1}-{\bm{R}}\bar{\mathbf{m}}+{\bm{R}}\bar{\mathbf{m}},\ldots,{\bm{R}}\mathbf{m}_{m}-{\bm{R}}\bar{\mathbf{m}}+{\bm{R}}\bar{\mathbf{m}}\})$
$=f(T_{{\bm{R}}\bar{\mathbf{m}}}({\bm{R}}M))$
$=f({\bm{R}}M)$ translation invariance
$=f(M)$ equation 31
$=f(T_{\bar{\mathbf{m}}}(M))$ translation invariance
$=f(M^{\prime}),$
which means that for any set of $m\leq n$ $n$-dimensional vectors $M^{\prime}$, an SE(n) invariant function $f$ has $f({\bm{R}}M^{\prime})=f(M^{\prime})$. Together with translation and rotation invariances of $f$, this implies that $f$ is E(n) invariant.
Remark 1.
If $f:\mathbb{R}^{k\times n}\rightarrow S,k\leq n$ is SE(n) invariant (equivalently E(n) invariant), and $g:S^{m}\rightarrow P$, then $g(f(M_{1}),\ldots f(M_{m})),$ with $M_{i}\in\mathbb{R}^{k\times n}$ is E(n) invariant.
The remark is made by applying any translation, rotation, and reflections to $M_{i}$ which leaves $f$ unchanged. And will be useful for our next proof.
B.1.5 Proposition 3 proof.
Proposition 3 proof.
Assume $n$ vectors $\mathbf{m}_{1},\ldots,\mathbf{m}_{n}$ in $\mathbb{R}^{3}$, where $n\geq 5$. Also, assume $f:\mathbb{R}^{m\times 3}\rightarrow S$ is an SE(3) invariant function, but not E(3) invariant. We know $m\geq 4$, by Lemma 2. Consider $f(\mathbf{m}_{1},\mathbf{m}_{2},\mathbf{m}_{3},\mathbf{m}_{4})=\operatorname{sign}(det([\mathbf{m}_{1}-\mathbf{m}_{4},\mathbf{m}_{2}-\mathbf{m}_{4},\mathbf{m}_{3}-\mathbf{m}_{4}]))$, and note that it is SE(3) invariant (rotations and translations leave the function unchanged, whereas reflections change its sign).
Remark 2.
Enantiomer configuration is a local property, which can only be determined by at least 4 vectors from its set of 5 atoms (4 residues and one central atom), and $f$ can be used to determine chirality. Throughout this proof, we refer to ”chirality” as the clockwise or counterclockwise atom arrangements when looking down from one vertex of a tetrahedron.
Let $M$ and $M^{\prime}$ be 4 atom subsets. Assume any triplet subset of $M$ has three linearly independent vectors. Then, $f(M)$ conveys no information on any other set $f(M^{\prime})$, unless $M=M^{\prime}$. Intuitively, choose any plane from $M$. The resulting determinant’s sign, which distinguishes chiral centers, will depend on what side of the plane we choose the 4th point. The fact that $f$ can be used to determine chirality stems from the equivalence between the classical chirality definition (using rotations) and determinant signs (see the beginning of Appendix B.1). All 4 sets of atoms need to be evaluated to have all potential enantiomer states that can happen in $n$ atom molecules.
We now observe that we do not strictly need to compute $f(\cdot)$ for all ordered sets of 4 atoms. For a set of 4 atoms, observe how their permutations may, at most, the sign:
$f(\mathbf{m}_{1},\mathbf{m}_{2},\mathbf{m}_{3},\mathbf{m}_{4})=f(\pi_{e}(\mathbf{m}_{1},\mathbf{m}_{2},\mathbf{m}_{3},\mathbf{m}_{4}))=-f(\pi_{o}(\mathbf{m}_{1},\mathbf{m}_{2},\mathbf{m}_{3},\mathbf{m}_{4})),$ (33)
where we denote by $\pi_{e}$ and $\pi_{o}$ even and odd permutation, respectively. Therefore, we need $\binom{n}{4}$ to compute all unordered sets of 4 atoms, giving $\mathcal{O}(n^{4})$.
Remark 3.
The $\binom{n}{4}$ features $f(\mathbf{m}_{i1},\mathbf{m}_{i2},\mathbf{m}_{i3},\mathbf{m}_{i4})$ need paired atom type information $(\mathbf{a}_{i1},\mathbf{a}_{i2},\mathbf{a}_{i3},\mathbf{a}_{i4})$ using a function that is order-variant on atoms, meaning.
$g(f(\mathbf{m}_{i1},\mathbf{m}_{i2},\mathbf{m}_{i3},\mathbf{m}_{i4}),(\mathbf{a}_{i1},\mathbf{a}_{i2},\mathbf{a}_{i3},\mathbf{a}_{i4}))\neq g(f(\mathbf{m}_{i1},\mathbf{m}_{i2},\mathbf{m}_{i3},\mathbf{m}_{i4}),(\pi(\mathbf{a}_{i1},\mathbf{a}_{i2},\mathbf{a}_{i3},\mathbf{a}_{i4}))),$ (34)
where $\pi$ is any permutation except the identity.
B.1.6 Cases where Proposition 3 applies
Consider the distributions $p_{1},\ldots,p_{n}$, with $p_{i}(\mathbf{m}_{i})=\mathcal{N}(\boldsymbol{0},\mathbf{I}),\forall i$ and random associated atom types $\{a_{i}\}$ (also drawn from an uninformative distribution). The task is transforming these distributions into the conformer distribution of $n$ atom molecules, some of which contain chiral centers. By the uninformative assumption, we need to compute the chirality of all sets of 5 atoms, since all sets have equal probability of being chiral. By remarks 1, 2 and 3, this scales prohibitively, in $\mathcal{O}(n^{4})$.
Chiral-aware, SE(3) invariance becomes intractable in point-cloud, diffusion models since the occurrence of reflections needs to be assessed in local neighborhoods (we are specifically interested in potential reflections of chiral centers, but all atoms may be part of chiral centers equally likely). For diffusion methods that define uninformative and identically distributed vector positions at $t=1$, it implies that all sets of 4 atoms may be part of tetrahedrons at $t=0$.
Examples of E(3) invariant diffusion models operating on point-clouds are (Hoogeboom et al., 2022; Peng et al., 2023; Vignac et al., 2023b; Xu et al., 2023; Wu et al., 2022). All these methods implement NN $\phi$ activating on sets of $\{f(M_{1}),\ldots,f(M_{n})\}$, where $M_{1}\in\mathbb{R}^{2\times 3}$, and $f$ are Euclidean norms $f(M_{i})=||M_{i,1}-M_{i,2}||_{2}$. They are E(3) invariant by Remark 1, and can only be made chiral-aware using $\mathcal{O}(n^{4})$ scaling, by Proposition 3.
B.2 Experiment details
@@
Appendix C Field parameters